# Eyes Can Deceive: Benchmarking Counterfactual Reasoning Capabilities of Multimodal Large Language Models

### v1

@@ 3.3. Dataset Construction
Annotation Pipeline. We ask annotators to perform counterfactual question annotations, which follows these steps: Given an image, annotators first select an appropriate question type. Images with no suitable question types will be excluded. Following the selected question type, annotators must initially craft an basic question to assess the models’ basic visual comprehension ability, then modify certain conditions within it to create a counterfactual question. To further enhance the data quality of CFMM, we follow a strict data quality control protocols. Specifically, each question is verified by at least one of the paper’s authors. We primarily examine the following two key points :
@@
(1) Information Leakage refers to the situation that answer to a question appearing in the ”If” clause, for example, ”If I painted this bus blue, what color wouold it be?” In this case, the model can derive answers directly from the text without paying attention to the image content.
@@
(2) The ambiguity of the answer refers to the inability to derive a precise answer to a question from the image. For example, it is not appropriate to ask about the number of elephants when only a small part of an elephant’s body is visible. We manually modified or removed all annotations with the above issues.
3.4. Evaluation Strategy
@@
Choice Extraction. Given the challenge of quantitatively evaluating open-ended answers, we structure the answers in the form of common multiple-choice questions. Specifically, each {basic question, counterfactual question} pair has two candidate answers. However, the existing instruction-following capabilities of MLLMs are limited. Even if explicitly asked to answer ”A” or ”B”, the model still produces some free-form text. Furthermore, we also observe that many MLLMs exhibit a preference for certain options, tending to favor those presented earlier. We adopt the same answer ranking strategy as SEED-Bench-2(Li et al., 2023a) to extract the model’s selection. Specifically, given a question, we compute the generation loss for each option and select the minimum loss as the model’s predicted result. This strategy eliminate the influence of option preferences and instructions, allowing for a more accurate reflection of the model’s actual capabilities.
@@
Evaluation Metrics. We utilize the accuracy and accuracy+ proposed in MME(Fu et al., 2023a), where accuracy here indicates correct responses to basic or counterfactual questions, and accuracy+ indicates that both questions were answered correctly. The random accuracy rates for these two are 50% and 25%, respectively. It can be seen that accuracy+ is a stricter metric, reflecting the extent to which the model fully comprehends the image. Unlike MME, we directly utilize the sum of accuracy and accuracy+ across all questions as our metric. Therefore, the full score of $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ are all 600.
@@
Model Prompt Count Color Size Shape Direction Common $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ Qwen-VL 0-shot 84.77 43.10 31.03 83.71 44.34 31.22 55.38 54.36 17.44 70.27 52.25 27.03 63.92 46.20 17.09 70.91 48.48 21.82 1-shot 84.77 59.48 47.99 83.71 46.15 35.29 55.38 55.90 20.03 70.27 56.76 31.53 63.92 50.00 24.05 70.91 52.73 26.06 CoT 84.77 58.05 46.55 83.71 46.15 35.29 55.38 54.36 17.95 70.27 50.45 25.23 63.92 48.73 24.05 70.91 53.94 27.27 ChatTruth 0-shot 84.77 65.81 53.16 81.90 52.49 38.01 62.56 47.69 14.87 63.06 46.85 14.41 63.90 51.90 22.15 71.52 52.73 24.85 1-shot 84.77 65.80 52.59 81.90 50.23 35.29 62.56 50.26 15.90 63.06 47.75 16.22 63.90 49.37 21.52 71.52 52.73 24.24 CoT 84.77 64.37 51.44 81.90 49.77 35.29 62.56 49.74 15.38 63.06 46.85 15.32 63.90 53.16 24.67 71.52 51.52 23.64 LLaVA 0-shot 87.36 39.37 28.74 77.83 46.15 24.89 63.59 47.18 16.92 76.58 59.46 38.74 63.29 50.00 18.35 75.15 53.94 31.52 1-shot 87.36 43.97 33.33 77.83 52.04 31.22 63.59 50.26 19.49 76.58 62.16 42.34 63.29 48.10 16.46 75.15 59.39 35.15 CoT 87.36 43.10 34.20 77.83 49.77 29.41 63.59 47.69 18.46 76.58 63.06 43.24 63.29 49.37 19.62 75.15 56.97 33.94 InfMLLM 0-shot 90.80 54.60 46.84 84.16 51.13 36.20 68.72 44.10 18.46 81.08 49.55 35.14 69.62 44.30 18.99 78.18 58.79 36.97 1-shot 90.80 50.57 42.82 84.16 52.94 38.46 68.72 43.08 16.41 81.08 54.05 39.64 69.62 43.04 18.99 78.18 56.36 34.55 CoT 90.80 48.85 41.09 84.16 51.58 37.10 68.72 45.13 18.46 81.08 54.95 40.54 69.62 44.30 19.62 78.18 57.58 36.36 InstructBLIP 0-shot 85.06 60.63 46.84 78.73 51.13 30.77 56.41 46.15 13.33 74.77 57.66 35.14 55.06 50.00 14.56 70.91 61.82 33.33 1-shot 85.06 55.17 45.40 78.73 56.56 44.80 56.41 44.62 18.46 74.77 50.45 30.63 55.06 45.57 24.69 70.91 49.70 27.88 CoT 85.06 52.30 42.82 78.73 56.11 42.53 56.41 46.15 20.00 74.77 51.35 32.43 55.06 48.73 22.15 70.91 50.91 27.27 BLIP2 0-shot 73.56 46.26 24.71 74.21 45.70 24.43 56.92 62.05 23.59 72.07 45.05 19.82 50.00 55.70 12.03 73.33 58.18 33.33 1-shot 73.56 47.99 28.45 74.21 52.04 33.48 56.92 44.10 15.90 72.07 47.75 27.03 50.00 55.06 24.05 73.33 46.67 27.27 CoT 73.56 53.45 35.63 74.21 51.13 36.20 56.92 46.67 18.97 72.07 46.85 25.23 50.00 45.57 16.46 73.33 50.30 32.73 Monkey 0-shot 86.78 54.31 43.39 85.07 49.32 35.29 54.36 55.38 17.44 64.86 50.45 18.02 60.76 50.63 18.35 70.91 47.27 19.39 1-shot 86.78 52.87 41.95 85.07 49.77 36.20 54.36 52.82 14.36 64.86 50.45 18.02 60.76 44.94 13.92 70.91 47.88 19.39 CoT 86.78 53.16 42.24 85.07 51.58 37.56 54.36 54.87 14.87 64.86 51.35 18.92 60.76 48.73 15.82 70.91 47.27 18.79
@@
Table 2. Performance of prevalent MLLMs on six tasks within our proposed CFMM. Here, $\text{acc}_{\text{b}}$ represents the accuracy for correctly answering basic questions, $\text{acc}_{\text{c}}$ denotes the accuracy for correctly answering counterfactual questions, and $\text{acc}_{\text{+}}$ signifies the accuracy for both types of questions. We highlight the best and second best results with red and blue, respectively.
3.5. Dataset Statistics
Overall, CFMM comprises 1200 images, each accompanied by one basic question and one counterfactual question. Figure LABEL:categories provides a detailed breakdown of the number of questions within each type. Note that despite our requirement for each annotator to provide balanced annotations for every question type, there still exists some imbalance among them. Specifically, ”count” are easier to formulate as counterfactuals, thus constituting a larger proportion of questions. In contrary, ”if” conditions in ”shape” often results in information leakage, rendering them more challenge to annotate. The distributions of the length of questions are shown in figures LABEL:bq and  LABEL:cq. Counterfactual questions necessitate the addition of ”if” conditions to basic questions, resulting in an average length of 18.39 words—much longer than that of basic questions, which is 9.47 words. Figure LABEL:wc display the word cloud of answers. For answers, we aim for simplicity and clarity, hence their average length is only 1.17 words.
3.6. Dataset Analysis
In Table 1, we list all datasets relevant to counterfactual reasoning. It can be observed that existing datasets primarily focus on Text QA, with relatively fewer explorations into VQA. The dataset most similar to ours is C-VQA. However, C-VQA is limited by the questions in VQA-v2, with most only involving quantity changes and boolean questions, while CFMM more comprehensively evaluates the counterfactual ability of the model from 6 different aspects. Moreover, we provide basic-counterfactual questions pair for each image, which can more intuitively reflect whether the model has truly engaged in counterfactual reasoning. Therefore, the CFMM we propose is a high-quality, comprehensive dataset for evaluating the counterfactual ability of the model.
4. Evaluation Results
In this section, we evaluate the performance of current state-of-the-art open source MLLMs on CFMM. We evaluated a total of 7 open source MLLMs, including Qwen-VL (Bai et al., 2023), ChatTruth, Monkey  (Li et al., 2023c), LLaVA (Liu et al., 2024), InfMLLM (Zhou et al., 2023), InstructBLIP (Dai et al., 2024) and BLIP2 (Li et al., 2023b). During the evaluation, we also tried two training-free prompt-tuning methods, 1-shot ICL and 1-shot CoT (Wei et al., 2022), to explore their impact on the counterfactual ability of MLLMs. Our evaluation results exhibit several interesting findings about these MLLMs and hope that these findings will be helpful for the development of the MLLMs community.
4.1. Main Results
The evaluation results of MLLMs across different question types are presented in Table 2 and an overview of the overall scores shown in Table 3. It can be seen that all MLLMs perform well on basic questions, achieving an accuracy of over 70%. For ”count” questions, InfMLLM even reaches an accuracy of 90.8%. By comparison, when evaluating counterfactual questions, a noticeable decline in model performance can be observed. All models perform close to or even below the level of random guessing (50%) in the 0-shot scenario. LlaVA exhibited the poorest performance in count-type questions, with an accuracy of only 39.37%. Overall, in terms of basic questions, InfMLLM ranks top-1, demonstrating the strongest multimodal understanding capability. InstructBLIP, due to the larger dataset scale during instruction tuning, possesses stronger command comprehension ability, hence achieving better performance on counterfactual questions.
@@ 4.1. Main Results
Qwen-VL 0-shot 428.96 288.73 145.63
1-shot 428.96 321.02 184.93
1-shot CoT 428.96 311.68 176.34
ChatTruth 0-shot 427.73 317.46 167.45
1-shot 427.74 316.13 165.76
1-shot CoT 427.73 315.41 165.75
LLaVA 0-shot 443.79 296.10 159.15
1-shot 443.79 315.92 177.99
1-shot CoT 443.79 309.97 178.87
InfMLLM 0-shot 472.57 302.47 192.59
1-shot 472.57 300.05 190.86
1-shot CoT 472.57 302.40 193.18
InstructBLIP 0-shot 420.95 327.39 173.97
1-shot 420.95 302.07 191.85
1-shot CoT 420.95 305.56 187.21
BLIP2 0-shot 400.10 312.94 137.92
1-shot 400.10 293.60 156.18
1-shot CoT 400.10 293.97 165.21
Monkey 0-shot 422.74 307.36 151.88
1-shot 422.74 298.73 143.85
1-shot CoT 422.74 306.96 148.2
Table 3. The overall performances of prevalent MLLMs on our proposed CFMM. $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ have the same meaning with those in Table 2. The “Total score” indicates the summation of scores achieved on six tasks. We highlight the best and second best results with red and blue, respectively.
4.2. Observations
Based on the above experimental results, we summarized four observations, hoping to be discussed by the community. We have also tried to analyze the possible causes and the directions available for further research:
@@ 4.2. Observations
All MLLMs we tested exhibited a significant performance drop when confronted with counterfactual questions, with an average decrease of 123.48 points in total score compared to $\text{acc}_{\text{b}}$. Notably, our analysis through $\text{acc}_{\text{+}}$ revealed a slight negative correlation coefficient in the models’ performance on basic questions and counterfactual questions, which indicates that when a model correctly answers a basic question, it is more likely to have committed an error on a counterfactual question. Likewise, the converse holds. This suggests a noteworthy phenomenon: even when a model provides correct answers to counterfactual questions, it may not have grasped the underlying counterfactual presuppositions. Instead, it might have simply relied on the wrong visual cues it obtained and chosen the “answer” corresponding to the basic question. Our CFMM counterfactual multimodal dataset has indeed posed formidable challenges to all existing state-of-the-art MLLMs. To date, no model has been able to effectively meet these challenges, and has demonstrated untrustworthy performance under counterfactual presuppositions.
Observation 2: The evaluated MLLMs perform best on numerical relation questions and worst on spatial positional relation questions.
@@
We evaluated MLLMs across 6 aspects. The experimental results indicate that MLLMs perform best in numerical relations, worst in spatial positional relations, and moderately in fundamental visual attributes and world knowledge. Existing MLLMs mostly utilize independent visual encoders like CLIP for extracting visual information, yet CLIP and similar encoders are trained for static short descriptions (Fu et al., 2023b; Radford et al., 2021). This makes CLIP and similar visual encoders most sensitive to object recognition, followed by sensitivity to object-level visual features, and less sensitive to spatial positional relations such as size and direction. As can be seen from the results, MLLMs exhibit a significant performance drop in counterfactual questions about counting and color, with $\text{acc}_{\text{c}}$ experiencing an average performance drop of over 30 points. When querying the visual information output by an independent visual encoder, object attributes are best obtained, but if the MLLM does not understand the counterfactual presuppositions properly, the queried visual information may instead lead to misleading answers. MLLMs perform well on world knowledge questions, which shows that 7B-level MLLMs have a certain amount of world knowledge and can eengage in appropriate common sense reasoning.
Figure 7. Qualitative results of ”Outputted a smooth but illogical CoT”
Observation 3: The introduction of 1-shot ICL provides a small improvement in model performance, but does not bring about a qualitative change.
@@
Experimental results show that introducing 1-shot ICL technique into the counterfactual reasoning task of MLLMs brings a visible performance improvement of more than 26 average total scores to Qwen-VL (Bai et al., 2023) and LLaVA (Liu et al., 2024). The introduction of 1-shot ICL appears to assist MLLM in better comprehending and leveraging contextual information, understanding the form of counterfactual reasoning tasks, thereby more accurately capturing the critical details of counterfactual presuppositions in the reasoning process and outputting them in the correct form. However, it is worth noting that although 1-shot ICL technology has brought some improvements, it has not fundamentally changed the MLLMs’ performance. 1-shot ICL has basically no effect on models such as ChatTruth that are further optimized on additional data, and there is a visible performance reduction on smaller scale models such as BLIP2 (Li et al., 2023b). This suggests that the improvement in enhancing the overall reasoning ability of the MLLMs is still limited. Previous studies have also shown that the ability of few-shot ICL improves with the increase in model’s parameter scale (Wei et al., 2023), which also explains one of the reasons why 1-shot ICL not perform well on the MLLMs we evaluated. Further explorations of the effectiveness of few-shot ICL in counterfactual reasoning on open-source MLLMs are also worth considering.
Observation 4: 1-shot CoT does not help much on MLLMs at the 7B level, and may even bring about a performance decline.
@@
Our experiments show that CoT does not significantly improve the model’s counterfactual reasoning ability for 7B-level MLLMs, and may even lead to performance degradation in some cases. In Wei et al.’s study (Wei et al., 2022), only models of a certain scale, such as exceeding 100B parameters, demonstrated advanced performance when combined with CoT strategies. CoT forces models into reasoning, yet 7B-level models often lack sufficient foundational knowledge and reasoning capabilities, making it unlikely for CoT to generate correct reasoning steps and potentially leading to issues like AI hallucinations. In Fig 7, we present an example of the output produced by Qwen-VL 7B when applying CoT. It can be seen that the model outputs fluent but illogical CoT, resulting in incorrect results. Ben et al. (Prystawski et al., 2024) also mentioned in their research that when using CoT, the model’s training data should be task-relevant and interconnected. CoT demonstrates excellent performance when the training data of large models exhibit local clusters of variables, yet the data for visual counterfactual reasoning is absent in MLLMs’ training data. The widespread application and deployment of MLLMs in industries and terminals necessitate that their scale cannot be too large. Meanwhile, the application of CoT will increase the regression length of the model, requiring more computational resources. Therefore, it is necessary to explore how CoT can be effectively implemented in smaller-scale MLLMs to reduce the cost of practical applications, enabling MLLMs to be more widely utilized.
5. Conclusion

### v3

@@ 3.3 Dataset Construction
Manually question annotation. We ask annotators to perform counterfactual question annotations, which follows these steps: Given an image, annotators first select an appropriate question type. Images with no suitable question types will be excluded. Following the selected question type, annotators must initially craft an basic question to assess the models’ basic visual comprehension ability, then modify certain conditions within it to create a counterfactual question. To further enhance the data quality of CFMM, we follow a strict data quality control protocols. Specifically, each question is verified by at least one of the paper’s authors. We primarily examine the following two key points:
  1. (1)
@@
    Information Leakage refers to the situation that answer to a question appearing in the “If” clause, for example, “If I painted this bus blue, what color wouold it be?” In this case, the model can derive answers directly from the text without paying attention to the image content.
  2. (2)
@@
    The ambiguity of the answer refers to the inability to derive a precise answer to a question from the image. For example, it is not appropriate to ask about the number of elephants when only a small part of an elephant’s body is visible. We manually modified or removed all annotations with the above issues.
Automated question expansion. We expanded our dataset based on 1,200 manually annotated question pairs. Leveraging GPT-4’s powerful multimodal understanding and reasoning capabilities, we prompted GPT to creatively generate new content by mimicking our manually annotated data. Specifically, we provided GPT with the original image, a basic question, a counterfactual question, and their respective answers. We then instructed GPT to modify the counterfactual condition while retaining the core question based on the image, thereby generating three new counterfactual questions. After this process, we ultimately obtained 6,000 questions. The prompts used for each category varied slightly, with the detailed prompts provided in the supplementary material. The newly generated questions were required to align with the image information and logical consistency, ensuring that the answers remain as concise and definitive as possible. We also incorporated manual verification to guarantee question quality in this stage.
3.4 Evaluation Strategy
@@
Choice Extraction. Given the challenge of quantitatively evaluating open-ended answers, we structure the answers in the form of common multiple-choice questions with two candidate answers. Considering that the existing instruction-following capabilities of MLLMs are limited. Even if explicitly asked to answer “A” or “B”, the models may still produce some free-form text. Furthermore, we also observe that many MLLMs exhibit a preference for certain options, tending to favor those presented earlier. We adopt the same answer ranking strategy as SEED-Bench-2(Li et al. 2023a) to extract the model’s selection. Specifically, given a question, we compute the generation loss for each option and select the minimum loss as the model’s predicted result. This strategy eliminate the influence of option preferences and instructions, allowing for a more accurate reflection of the model’s actual capabilities.
@@
Evaluation Metrics. We utilize the accuracy and accuracy+ proposed in MME(Fu et al. 2023a), where accuracy here represent correct responses to basic or counterfactual questions, and the accuracy+ indicates that both questions were answered correctly. The random accuracy rates for these two metrics are 50% and 25%, respectively. It can be seen that accuracy+ is a stricter metric, reflecting the extent to which the model fully comprehends the image. Unlike MME, we directly utilize the sum of accuracy and accuracy+ across all questions as our metric. Therefore, the full score of $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ are all 600.
@@
Model Prompt Count Color Size Shape Direction Common $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ Qwen-VL 0-shot 86.78 50.65 42.24 89.14 74.43 67.99 69.23 52.05 34.23 90.09 51.35 42.57 69.62 54.43 33.86 78.79 58.18 41.97 1-shot ICL 86.49 64.44 55.24 87.33 76.47 70.25 69.23 61.92 41.41 89.19 57.43 48.65 71.52 55.70 34.34 78.79 69.70 52.12 1-shot CoT 89.37 62.14 53.02 90.95 76.13 69.80 69.35 60.64 40.77 85.59 55.18 46.62 67.09 57.28 35.60 77.81 67.27 49.85 ChatTruth 0-shot 87.36 70.33 61.49 92.31 76.70 70.93 75.38 55.77 41.53 82.88 58.78 44.59 74.68 49.53 35.60 76.97 66.82 48.03 1-shot ICL 86.21 69.54 60.70 92.31 76.58 70.81 73.84 55.51 41.41 82.88 59.46 45.27 71.52 50.63 37.18 77.58 66.36 47.58 1-shot CoT 87.65 68.61 59.70 90.05 76.36 70.59 74.35 53.72 39.87 81.98 59.23 45.50 70.88 50.00 36.87 73.33 66.67 47.88 LLaVA 0-shot 87.36 50.65 41.88 90.05 71.27 64.03 59.49 53.85 30.64 74.77 69.59 48.65 66.46 56.33 33.23 75.15 64.85 43.94 1-shot ICL 85.64 53.66 45.19 90.05 73.64 66.63 58.98 54.23 30.90 74.77 70.50 50.23 65.19 55.70 33.70 76.97 65.91 45.15 1-shot CoT 81.04 53.66 45.04 87.34 71.95 64.93 56.41 56.15 31.67 76.57 71.17 50.45 63.3 54.59 32.75 73.94 64.70 43.79 InfMLLM 0-shot 91.09 63.94 57.61 93.21 72.40 67.42 66.15 55.38 36.28 78.38 66.44 50.00 74.68 55.06 36.08 82.42 68.03 52.88 1-shot ICL 92.53 61.42 54.89 89.14 72.29 67.31 66.12 55.13 35.90 79.28 65.32 48.87 75.31 54.91 36.71 84.24 67.12 52.42 1-shot CoT 89.08 60.13 53.52 91.85 72.29 67.31 70.77 56.15 36.79 77.48 67.12 50.90 74.05 54.75 36.71 81.21 68.18 53.33 InstructBLIP 0-shot 84.77 67.39 56.61 90.50 72.74 65.72 58.97 56.79 31.28 79.28 55.86 41.89 56.33 53.64 29.59 80.61 61.52 48.79 1-shot ICL 86.78 51.44 42.67 90.5 52.38 46.83 58.66 48.46 28.72 78.48 38.96 28.60 55.7 47.63 26.27 78.39 38.94 28.94 1-shot CoT 80.75 48.20 40.37 85.98 54.64 48.87 59.02 48.85 28.21 81.08 38.06 28.83 50.63 49.53 25.63 78.72 39.24 27.88 BLIP2 0-shot 74.43 53.66 37.72 81.45 69.91 57.13 66.67 65.77 42.95 73.87 52.25 32.21 53.16 52.06 23.26 79.39 61.67 47.58 1-shot ICL 74.71 42.74 28.74 81.9 54.64 45.02 63.59 47.82 28.97 72.97 35.14 20.95 51.26 45.25 22.15 78.18 38.48 29.09 1-shot CoT 72.99 49.93 34.77 81.01 52.94 44.34 66.67 49.62 30.64 71.17 38.29 24.32 56.96 41.61 20.89 81.21 40.61 30.76 Monkey 0-shot 88.79 61.14 53.30 92.76 73.42 68.10 69.74 58.59 39.23 88.29 52.70 42.57 74.22 50.32 34.81 84.33 64.55 52.42 1-shot ICL 89.65 60.06 52.30 92.31 73.30 68.10 69.23 57.44 38.85 88.29 49.55 39.86 75.32 49.21 33.23 84.85 64.55 52.73 1-shot CoT 87.64 59.77 51.87 91.73 73.19 67.87 68.71 57.44 38.97 89.75 51.80 42.12 76.59 49.37 33.54 83.98 65.45 53.64
@@
Table 2: Performance of prevalent MLLMs on six tasks within our proposed CFMM. Here, $\text{acc}_{\text{b}}$ represents the accuracy for correctly answering basic questions, $\text{acc}_{\text{c}}$ denotes the accuracy for correctly answering counterfactual questions, and $\text{acc}_{\text{+}}$ signifies the accuracy for both types of questions. We highlight the best and second best results with red and blue, respectively.
@@
Qwen-VL 0-shot 483.65 341.10 262.86
1-shot ICL 482.55 385.66 302.01
1-shot CoT 480.16 378.64 295.65
ChatTruth 0-shot 489.58 377.93 302.17
1-shot ICL 484.34 378.08 302.95
1-shot CoT 478.23 374.59 300.41
LLaVA 0-shot 453.27 366.53 262.37
1-shot ICL 451.60 373.64 271.79
1-shot CoT 438.60 372.22 268.63
InfMLLM 0-shot 485.94 381.25 300.27
1-shot ICL 486.62 376.18 296.10
1-shot CoT 484.44 378.61 298.57
InstructBLIP 0-shot 450.46 367.93 273.88
1-shot ICL 448.51 277.80 202.03
1-shot CoT 436.18 278.52 199.79
BLIP2 0-shot 428.97 355.32 240.83
1-shot ICL 422.61 264.08 174.92
1-shot CoT 430.01 272.99 185.72
Monkey 0-shot 498.13 360.72 290.43
1-shot ICL 499.65 354.11 285.07
1-shot CoT 498.40 357.02 288.01
Table 3: The overall performances of prevalent MLLMs on CFMM. $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ have the same meaning with those in Table 2. The “Total score” indicates the summation of scores achieved on six tasks. We highlight the best and second best results with red and blue, respectively.
3.5 Dataset Analysis
In Table 1, we list all datasets relevant to counterfactual reasoning. It can be observed that existing datasets primarily focus on Text QA, with relatively fewer explorations into VQA. The dataset most similar to ours is C-VQA. However, the types of questions and answers in C-VQA are limited, and we also superior in terms of scale, CFMM can more comprehensively evaluates the counterfactual ability of the model from different aspects. Moreover, we provide basic-counterfactual questions set for each image, which can more intuitively reflect whether the model has truly engaged in counterfactual reasoning. Therefore, the CFMM we propose is a high-quality, comprehensive dataset for evaluating the counterfactual ability of the model.
4 Evaluation Results
In this section, we evaluate the performance of current state-of-the-art open source MLLMs on CFMM. We evaluated a total of 7 open source MLLMs, including Qwen-VL (Bai et al. 2023), ChatTruth, Monkey  (Li et al. 2023e), LLaVA (Liu et al. 2024), InfMLLM (Zhou et al. 2023), InstructBLIP (Dai et al. 2024) and BLIP2 (Li et al. 2023d). During the evaluation, we also tried two training-free prompt-tuning methods, 1-shot ICL and 1-shot CoT (Wei et al. 2022), to explore their impact on the counterfactual ability of MLLMs. The CoT prompts are based on the proven approach outlined in ScienceQA (Lu et al. 2022), utilizing manually annotated CoT processes as in-context learning examples. Our evaluation results exhibit several interesting findings about these MLLMs and hope that these findings will be helpful for the development of the MLLMs community.
4.1 Main Results
Our CFMM has posed formidable challenges to all existing state-of-the-art MLLMs. To date, no model has been able to effectively meet these challenges, and their performance under counterfactual presuppositions has proven unreliable. The evaluation results of MLLMs across different question types and overall scores are presented in Table 2 and Table 3 respectively. It can be seen that all MLLMs perform well on basic questions, achieving an accuracy of around 80%. For “color” questions, InfMLLM even reaches an accuracy of 93.21%. By comparison, a noticeable decline can be observed when evaluating counterfactual questions. All models perform close to the level of random guessing (50%) in the 0-shot scenario. LLaVA exhibited the largest performance drop on counting problems, with an accuracy of only 50.63% on counterfactuals, but an accuracy of 87.36% on basic problems. Overall, Monkey performs best on basic questions, demonstrating the strongest multimodal understanding capability. InfMLLM, due to the introduction of extra VQA datasets while training, possesses stronger query comprehension ability, hence achieving better performance on counterfactual questions.
4.2 Observations
Based on the above experimental results, we summarized four key observations, analyzed their potential causes, and suggest directions for future research.
@@ 4.2 Observations
All MLLMs exhibited a significant performance drop when confronted with counterfactual questions, with an average decrease of 105.75 points in total score compared to $\text{acc}_{\text{b}}$. Notably, our analysis through $\text{acc}_{\text{+}}$ revealed a slight negative correlation coefficient in the models’ performance on basic questions and counterfactual questions, which indicates that when a model correctly answers a basic question, it is more likely to have committed an error on a counterfactual question. Likewise, the converse holds. This suggests a noteworthy phenomenon: even when a model provides correct answers to counterfactual questions, it may not have grasped the underlying counterfactual presuppositions. Instead, it might have simply relied on the wrong visual cues it obtained and chosen the “answer” it thought corresponding to the basic question. In most cases, visual information is accurate, and models rely on this information while ignoring counterfactual presuppositions, leading to a significant performance drop on $\text{acc}_{\text{c}}$.
Observation 2: The evaluated MLLMs are most sensitive to single-object attributes and worst sensitive to relative relation attributes.
@@
Table2 indicates that MLLMs perform best on basic questions involving single-object attributes such as count and color, but perform worst on questions involving multiple-object attributes related to spatial positional relations, such as size and direction. Existing MLLMs mostly utilize independent visual encoders like CLIP for extracting visual information, yet CLIP and similar encoders are trained for static short descriptions (Fu et al. 2023b; Radford et al. 2021). This makes CLIP and similar visual encoders most sensitive to object recognition and object-level visual features, but less sensitive to spatial positional relations. As can be seen from the results, MLLMs exhibit a significant performance drop in counterfactual questions about counting and color, with $\text{acc}_{\text{c}}$ dropping by over 20 points. When querying the visual information output by an independent visual encoder, object attributes are easiest obtained, but if the MLLM does not understand the counterfactual presuppositions properly, the queried visual information may instead lead to misleading answers.
Figure 3: Qualitative results of “Outputted a smooth but illogical CoT”.
Observation 3: The improvement brought by 1-shot ICL is inconsistent.
Experimental results show that introducing 1-shot ICL into the counterfactual reasoning task significantly improved Qwen-VL’s total score by more than 40 points.
@@
The introduction of 1-shot ICL appears to help MLLM better understand and utilize contextual information, enabling it to more accurately capture key details of counterfactual presuppositions in the reasoning process. However, it is worth noting that while 1-shot ICL has brought some improvements, it is not effective for all MLLMs. For instance, 1-shot ICL has basically no effect on models such as ChatTruth and InfMLLM that are optimized with additional data, and the performance noticeably drops in models that have undergone additional instruction tuning, such as InstructBLIP, or in smaller scale models like BLIP2. Previous studies have shown that the few-shot ICL performance improves as the model’s parameter scale increases (Wei et al. 2023), which may explain why 1-shot ICL did not perform well on the MLLMs we evaluated. Further explorations of the effectiveness of few-shot ICL in counterfactual reasoning on open-source MLLMs are also worth considering.
Observation 4: 1-shot CoT does not help much on MLLMs at 7B-level.
@@
Similar to 1-shot ICL, 1-shot CoT does not significantly enhance the counterfactual reasoning ability of 7B-level MLLMs and may even lead to performance drop in some cases. In Wei et al.’s study (Wei et al. 2022), only models of a certain scale, specifically those exceeding 100B parameters, demonstrated advanced performance when combined with CoT strategies. CoT forces models to engage in reasoning, yet 7B-level models often lack sufficient foundational knowledge and reasoning capabilities, making it unlikely for CoT to generate correct reasoning steps and potentially leading to issues like AI hallucinations. For instance, Figure 3 shows an output example produced by Qwen-VL 7B when using CoT. It can be seen that the model outputs fluent but illogical CoT, resulting in incorrect results. Ben et al. (Prystawski, Li, and Goodman 2024) also mentioned in their research that when using CoT, the model’s training data should be task-relevant and interconnected. CoT demonstrates excellent performance when the training data of large models exhibit local clusters of variables. However, visual counterfactual reasoning data is absent in the training datasets of MLLMs. Therefore, it is necessary to explore how CoT can be effectively implemented in smaller-scale MLLMs to reduce the cost of practical applications, enabling MLLMs to be more widely utilized.
5 Conclusion